OpenStudy (anonymous):

How do you determine if a function is even, odd or neither? y=3x^2+4

OpenStudy (anonymous):

Even function: f(x) = f(-x). Odd function: -f(x) = f(-x)

OpenStudy (anonymous):

Your function is Even. Because (-x)^2 = (x)^2

OpenStudy (anonymous):

\[f(-x)=\cases{\begin{array}{cl} f(x) & \text{if } f \text{ is even}\\ -f(x) & \text{if } f \text{ is odd}\end{array}}\] Otherwise neither.

OpenStudy (anonymous):

"Because (-x)^2 = (x)^2" how did you figure that out?

OpenStudy (anonymous):

Every negative number when is "^2" turns into a positive number